OpenStudy (anonymous):

Part 1 Is the following equation true? Part 2 Use complete sentences to explain the properties used in making your decision. 4 log 3 x + log 3 y – 2 log 3 z = log 3 .

OpenStudy (anonymous):

I think it's true because they are all log 3 right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it depends on what \(x,y,z\) are

OpenStudy (anonymous):

\[4 \log_3 (x) + \log_3(y) – 2 \log_3(z) = \log 3 \] \[\log_3(\frac{x^4y}{z^2})=\log(3)\] depends on the variables

OpenStudy (anonymous):

mmmkkk so then it is not a true function because those are different?